OpenStudy (anonymous):

Please Help!!!! Medal and Fan!!!!! @pawtpie Cards with the numbers 1 through 50 are placed in a box and thoroughly shuffled. What is the probability of randomly selecting a card with a number less than 10 or a multiple of 12?

OpenStudy (anonymous):

How many numbers are less than 10?

OpenStudy (anonymous):

And >= 1

OpenStudy (anonymous):

\[A. \frac{ 36 }{ 2500 }\] B. \[\frac{ 13 }{ 100 }\] C. \[\frac{ 13 }{ 50 }\] D. \[\frac{ 7 }{ 25 }\]

OpenStudy (anonymous):

How many numbers are less than 10 but greater than or equal to 1

OpenStudy (anonymous):

9

OpenStudy (anonymous):

What are the multiples of 12 below 15

OpenStudy (anonymous):

2*6 3*4

OpenStudy (anonymous):

12, 24, 36, and 48

OpenStudy (anonymous):

And there's 4 of these numbers

OpenStudy (anonymous):

So in total that's 13 out of the 50

OpenStudy (anonymous):

or 13/50

OpenStudy (anonymous):

I have another one A bag contains 5 red cubes and 7 white cubes. Celinda will select one cube, put it in her pocket, and then select a second cube. Which expression given the probability that she will select two red cubes?